OpenStudy (anonymous):

q + 12 - 2(q - 22) > 0

OpenStudy (anonymous):

@myininaya

OpenStudy (adamaero):

multiply that 2 out

myininaya (myininaya):

He is referring the distributive property. you have a(b+c) above a(b+c)=ab+ac by the distributive property

myininaya (myininaya):

so what is -2(q-22) equal to using that property?

OpenStudy (anonymous):

dont get it ...

myininaya (myininaya):

Example: 3(10-x)=3(10)-3(x)=30-3x By the distributive property.

myininaya (myininaya):

So you try multiplying -2 and (q-22)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

now what

myininaya (myininaya):

I would combine like terms after that

myininaya (myininaya):

what did you get when multiplying -2 and (q-22)

OpenStudy (anonymous):

alright, i get it now thannk you

OpenStudy (anonymous):

@myininaya im not sure if i did it right, can you explain further ?

myininaya (myininaya):

Show me what you did.

myininaya (myininaya):

-2(q-22) =-2(q)-(-2)(22) =-2q+44

myininaya (myininaya):

\[q+12-2(q-22) \ge 0 \\ q+12-2q+44 \ge 0\]

myininaya (myininaya):

now combine like terms on the left hand side

OpenStudy (anonymous):

so its not negative ?

OpenStudy (anonymous):

alright

myininaya (myininaya):

-2(-22) is positive 44

myininaya (myininaya):

now combining like terms what is q-2q=? what is 12+44=?

OpenStudy (anonymous):

56

myininaya (myininaya):

yes 12+44=56 and q-2q=?

OpenStudy (anonymous):

it might sound stupid but im just catching up in all my work so this is why im on openstudy..

myininaya (myininaya):

that is fine and good \[q+12-2(q-22) \ge 0 \\ q+12-2q+44 \ge 0 \\ q-2q+56 \ge 0\]

OpenStudy (anonymous):

okay, thank you so much

myininaya (myininaya):

what is q-2q?

myininaya (myininaya):

1-2=-1 so q-2q=?

OpenStudy (anonymous):

??

myininaya (myininaya):

1q-2q=-1q

OpenStudy (anonymous):

okkayy

myininaya (myininaya):

\[q+12-2(q-22) \ge 0 \\ q+12-2q+44 \ge 0 \\ q-2q+56 \ge 0 \\ -q+56 \ge 0\]

myininaya (myininaya):

Only one last step or two if you choose to go the other way.

OpenStudy (anonymous):

Ive submitted my test already and i did the last step, so thank you so much

myininaya (myininaya):

Oh no. I didn't know I was helping you cheat. I really think in the future you should take your tests using your own knowledge even if behind. A test is to test one's on knowledge.

OpenStudy (anonymous):

this and one other question is really all i asked help with. it was only because i was two lessons behind but now im fine with my algebra lessons, everything else i did alone so it wasnt really cheating because you helped and explained it to me so thnk you for that